# ProxNF: Neural Field Proximal Training for High-Resolution 4D Dynamic Image Reconstruction

### v1

@@ III-B Adapted partition of unity neural fields
In this work, the regularization term $\mathcal{R}$ was chosen to be Tikhonov regularization on the on the spatial derivatives computed with finite difference and on the temporal derivatives computed with automatic differentiation of the learned partition of unity. This choice of regularization then allowed the proximal $\operatorname{Prox}_{\alpha,\mathcal{R}}^{\textnormal{NF}}$ to be computed using the alternating minimization algorithm outlined in Eq. (13), where the coefficient updates can be accomplished by solving a linear system while updates of the partition update can be updated using a stochastic gradient method.
@@ IV Numerical Studies
This work utilized two numerical studies to assess the proposed NF representation and proposed ProxNF method for spatiotemporal image reconstruction. These studies were inspired by dynamic contrast-enhanced (DCE) photoacoustic computed tomography (PACT) imaging of tumor perfusion in small animal models [41]. PACT is a hybrid biomedical imaging modality that combines benefits from both optical and ultrasound imaging[42]. Specifically, PACT combines the benefits of high soft tissue contrast from optical imaging and high resolution detection principles from ultrasound imaging to create high resolution maps of optical properties. PACT image formation utilizes a short laser pulse in the near infrared range to illuminate a biological tissue. The biological tissue then absorbs the optical energy and generates heat and a local increase in pressure, which is described as the photoacoustic effect. This pressure distribution then propagates as an acoustic wave where it can be measured outside the tissue using ultrasound detectors. Under the assumption of a non-lossy homogeneous acoustic medium, that is, a medium with homogeneous speed of sound and density and no acoustic attenuation, the relationship between the induced pressure distribution and the data can be modeled using the circular Radon transform in two spatial dimensions and the spherical Radon transform in three spatial dimensions [43, 44]. The numerical studies presented in this work focus on estimation of this spatially and temporally varying induced pressure distribution from sparse-view measurements collected from a circular aperture in two spatial dimensions and a cylindrical aperture in three spatial dimensions.
@@ IV-A Construction of the dynamic object
The dynamic objects used in these studies were spatiotemporal maps of the spatially and temporally varying induced pressure distribution within a numerical tumor perfusion model in two and three spatial dimensions. The anatomical structures of these objects were constructed utilizing an abdominal section extracted from an anatomically realistic whole-body murine numerical phantom (MOBY) [45, 46]. MOBY comprises detailed three-dimensional (3D) anatomical structures featuring tens of organs and models physiologically realistic cardiac and respiratory motions. A spherical lesion of diameter 7 mm was inserted in the abdominal region at a depth of 5 mm to mimic a subcutaneous tumor flank. To simulate DCE imaging of a tumor perfusion model, the concentration of an optical contrast agent in the lesion and blood-filled organs was computed using a two compartment Toft’s model [47, 48, 49]. In this model, the contrast agent was introduced after 90 seconds and injected over a 30 second window following a slow bolus pattern. Optical properties (optical absorption coefficient and scattering coefficient) corresponding to a 800 nm laser pulse wavelength were then assigned to each tissue type based on physiologically plausible chromophore concentrations (oxygenated hemoglobin, deoxygenated hemoglobin, water, fat, contrast agent)[50]. Contrast agent optical properties were chosen to mimic those of indocyanine green (ICG) [51], in which case the absorption spectrum has a peak near the chosen wavelength.
The diffusion approximation[52] was used to simulate photon transport at each imaging frame. The incident fluence was assumed constant on the surface of the object. The fluence distribution within the object was obtained by solving the partial differential equation stemming from the diffusion approximation using the open-source finite element library FEniCS [53], thus yielding a volumetric map of dynamic induced pressure distribution at each imaging frame. For the experiments in two spatial dimensions a slice was extracted from the 3D induced pressure phantom.
Figure 1: Numerical dynamic mouse abdominal phantom extract from mouse whole body phantom (MOBY) and shown at beginning and and of inspiration during a 5 second breathing period. The phantom is comprised of detailed three-dimensional (3D) anatomical structures featuring several of organs and models physiologically realistic cardiac and respiratory motions. A spherical lesion of diameter 7 mm was inserted in the abdominal region at a depth of 5 mm to mimic a subcutaneous tumor flank and displayed in dark red.
@@ IV-B1 Dynamic imaging system for two spatial dimensions
Figure 2 illustrates the virtual imaging system used in the numerical study with two spatial dimensions. The measurements geometry was circle of radius $R$ (shown in green), along which two groups of 5 idealized point-like ultrasound transducers are distributed. The sensors in each group were spaced $1^{\circ}$ apart on the transducer arc. Each group of sensors was distributed antipodal from one another. The object was placed at the center of the imaging system and fully contained within a square region of size $L$ (field of view). At each imaging frame, data were collected simultaneously by each sensor corresponding to the integral along $I$ concentric arcs. Sensors were rotated by an angle $\Delta\Theta$ after each imaging frame. The dynamic changes of the object were assumed to be slow enough such that the object was not expected to vary significantly during the acquisition of each imaging frame (quasi-static approximation). This assumption is often valid for several dynamic PACT applications since the duration of the light excitation pulse at each imaging frame is in the range of a few nanoseconds while the dynamic changes in the object have a time-scale in the range of milliseconds to seconds.
Under these assumptions, the imaging operator in two spatial dimensions can then be described by the circular Radon transform (CRT). For a sensor location $r_{0}^{(k)}\in\mathbb{R}^{2}$, measurement radius $\ell$, and imaging frame $t_{k}$, the CRT can be expressed as the integral along a circular arc
[…]
The superscript $k$ in $r_{0}^{(k)}$ is used to highlight the fact that the location $r_{0}^{(k)}$ of the sensor changes at each imaging frame. The discrete-to-discrete counterpart of the CRT imaging operator was implemented using the AirToolsII MatLab toolbox [54] and stored as a compressed sparse column matrix[55]. The imaging systems parameters are displayed in Table I.
Figure 2: Illustration of the virtual system for dynamic imaging using a CRT model. At each imaging frame, data are collected by idealized point-like transducers (shown in blue). Transducers rotate of an angle $\Delta\Theta$ between frames.
@@ IV-B2 Dynamic imaging system for three spatial dimensions
The object was placed at the center of the imaging system and fully contained within a rectangular prism region of side length $L$ in the $x$ and $y$ direction and height $H$. At each imaging frame, data were collected at each sensor ($I$ measurements per sensor). Sensors rotated an angle $\Delta\Theta$ along the z-axis of the cylindrical geometry at each frame. Again the dynamic changes in the object were assumed to be slow enough such that the object is approximately constant at each imaging frame.
@@ IV-C Study Design
The first numerical study solved the embedding problem and assessed the proposed NF architecture for representing the dynamic object described in Section IV-A. For the case involving two spatial dimensions, the proposed NF architecture was compared to the original NF architecture with polynomial basis functions [13] and a low-rank (semiseparable) decomposition of the dynamic object. The proposed POUnet architecture used a network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ with 4 hidden layers of width 140 nodes and one output layer of width $P=10$ partitions. Sinusoidal and softmax activation functions were employed in the hidden and output layers, respectively. The set of basis function $\mathbf{B}$ consisted of $L=186^{2}=34596$ pixel indicator functions. This resulted in a representation with $427,020$ parameters. The original POUnet architecture used a network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ with 5 hidden layers of width 280 nodes and output layer of width $P=100$ partitions. Third order space-time separable polynomials were employed as basis functions ($L=40$). This resulted in a representation with $425,800$ parameters. These two NF representations were compared to a semiseparable approximation of rank $r=12$, SS(12), which has $430,704$ parameters. Comparatively, storing the dynamic object as a collection of spatiotemporal voxels requires $\sim 20$ million parameters. Therefore these representations allow for a 40-fold reduction in the number of necessary parameters.
For the case involving three spatial dimensions, the proposed NF architecture was compared to a low-rank, semiseparable, decomposition of the dynamic image. The POUnet architecture again utilized the same network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ as the 2D case and a set of basis function $\mathbf{B}$ that contained $L=186^{2}\cdot 100=3,459,600$ voxel indicator functions. This resulted in a representation with $34,988,424$ parameters. The semiseparable approximation used a rank $r=10$, SS(10), which has $34,608,960$ parameters. These representations yield over two-order of magnitude reduction in the number of parameters compared to representing the object as a collection of spatiotemporal voxels values ($\sim 5$ billion parameters).
@@
The second numerical study explored dynamic image reconstruction in two (CRT imaging operator) and three (SRT imaging operator) spatial dimensions. In both cases, two spatiotemporal image reconstruction approaches were utilized. The first reconstruction technique utilized the proposed NF architecture and the proposed ProxNF method in Algorithm 1. The second reconstruction technique, which serve as a reference, was spatiotemporal image reconstruction with nuclear norm regularization (STIR-NN) [58]. In particular, for the case involving three space dimensions, the memory-efficient maximal-rank-constrained method in [11] was implemented. Measurement data were corrupted with i.i.d. additive Gaussian noise with zero mean and standard deviation $\sigma$ controlled by a specified relative noise level (RNL), defined as $\operatorname{RNL}=\sigma/\|\boldsymbol{d}\|_{\infty}$. In the two spatial dimensions case, the RNL was set to 0.04 and, in the three spatial dimensions case, the RNL was set to 0.01. Regularization weights for all reconstructions were chosen using Morozov’s discrepancy principle [59] in which the regularization weight was swept so that the standard deviation of the data-fidelity offset was approximately equal to the noise standard deviation. In the two spatial dimensions case CRT data were collected using 2 rotating sets of sensors (5 sensors per set) placed antipodally apart. In the three spatial dimensions case, SRT data were generated from 4 rotating sets of sensors placed 90 degree apart from each other.
@@
In these studies, image quality was quantified using relative root mean square error (RRMSE) and structural similarity index measure (SSIM) [60], evaluated both on the entire image and on a region of interest containing the lesion. Image quality was also assessed by analyzing the estimated time activity curve, which is the average signal over a volume of interest (the lesion).
@@ V Results
The proposed method, ProxNF, was implemented using PyTorch, an open-source machine learning framework [61], and executed on GPUs. The POUnet parameters $\boldsymbol{\eta}$ and $\boldsymbol{C}$ were initialized by performing an embedding of the time-averaged object for the representation studies and by solving a static image reconstruction problem on time-averaged measurements for the reconstruction studies.
Following initialization, the NFs in the representation study were trained using the alternating minimization approach in Eq. (13). Similarly after initialization, the NFs in the reconstruction study were trained with the proposed proximal splitting method in Algorithm 1. A reduction of the order of magnitude in the norm of the proximal gradient was used as stopping criterion. The batch size was $J=324$ and time frames were stochastically selected at each iteration. Tikhonov regularization was used to penalize the $L^{2}$ norm of the gradient of the spatiotemporal image. The regularized embedding problem at each proximal update was solved as follow. The linear systems corresponding to the updates of $\boldsymbol{C}$ were solved using the conjugate gradient method. The internal updates for $\boldsymbol{\eta}$ were accomplished using an Adam optimizer [24]. Specifically, 10,000 iterations with a fixed learning rate of $10^{-5}$ and batch size of 100,000 randomly sampled spatiotemporal points were used.
The spatiotemporal reconstructions in two spatial dimensions using the reference method (STIR-NN) were accomplished utilizing the Fast Iterative Shrinkage Algorithm (FISTA) [62] as implemented in UNLocBoX, a MATLAB toolbox for proximal-splitting methods [63]. The low rank semiseparable approximation of the image with two spatial dimensions was accomplished using SVD command in MATLAB and soft-thresholding. The spatiotemporal reconstructions in three spatial dimensions using the reference method (STIR-NN) were accomplished using the memory-efficient randomized algorithm in [11] with a maximum rank constraint of 32. The randomized singular value decomposition [64] was employed to compute the low rank decomposition and enforce the maximum rank constraint.
@@ V-A Study 1: Dynamic embedding problem
Figure 3: Study 1: Average lesion activity of true object in two spatial dimensions (blue), SS(12) approximation (green), original POUnet architecture (purple), and proposed POUnet architecture (red). The proposed POUnet architecture better captures the dynamic changes of the object compared to the original POUnet architecture in [13].
Figure 4: Study 1: Dynamic object in two spatial dimensions (top), Rank 12 semiseparable approximation SS($12$) (second row), neural field using original POUnet architecture (third row), and neural field using proposed POUnet architecture are shown at selected frames throughout time period. The dynamic range of the grayscale image is $041e-$ 04 to $0.06$. A logarithmic scale was used in the color map to increase visualization contrast at depth. An animation of each representation constructed in this study is available in the supplemental multimedia materials (Video 1).
Figure 5: Study 1: Average lesion activity of true object in three spatial dimensions (blue), the semi-separable rank 10 approximation (green), and the neural field representation (red). The semiseparable and neural field representation lead to an almost identical, smooth lesion activity curve that underestimates activity at later times.
Figure 6: Study 1: Dynamic object in three spatial dimensions (top), Rank 10 semiseparable approximation SS($10$) (second row), neural field representation (bottom row) with proposed POUnet architecture are shown at selected frames throughout time period. Paraview [65] was used for volume rendering. Dynamic range of the rendering is [0-0.06] using a linear scale. Both methods provide high fidelity representations but do slightly underestimate the tumor activity in later time periods. An animation of each representation constructed in this study is available in the supplemental multimedia materials (Video 2).
TABLE II: Study 1: Dynamic embedding problem. RRMSE, SSIM, ROI RRMSE, ROI SSIM, and LAC RRMSE achieved by the NF representations with original and proposed POUnet architectures and the rank $r$ semiseparable approximations SS($r$) in two (2D) and three (3D) spatial dimensions.
@@
The accuracy metrics—RRMSE and SSIM over the entire image, RRMSE and SSIM over the region of interest containing the inserted lesion, and the lesion activity curve RRMSE— achieved by each representation are reported in Table II. The lesion activity curves associated with these representations and with the object are shown in Fig. 3. Selected frames of these representations and of the object are shown in Fig. 4. The proposed NF architecture showed superior approximation properties (as quantified by the above accuracy metrics) compared to the original POUnet architecture in [13], and comparable accuracy compared to the semiseparable approximation with a similar number of parameters.
@@
The accuracy metrics for the proposed POUnet representation and the semiseparable approximation of rank 10 (SS(10)) are reported in Table II. The average lesion activity curves associated with these representations and with the object are shown in Fig. 5. Selected frames of these representations and the object are shown in Fig. 6. The proposed NF architecture achieved comparable accuracy metrics to those achieved by the semiseparable approximation with a similar number of parameters.

### v3

@@ III-C Adapted partition of unity neural fields
In this work, the regularization term $\mathcal{R}$ was chosen to be Tikhonov regularization on the spatial derivatives computed with finite difference and on the temporal derivatives computed with automatic differentiation of the learned partition of unity. This choice of regularization then allowed the proximal $\operatorname{Prox}_{\alpha,\mathcal{R}}^{\textnormal{NF}}$ to be computed using the alternating minimization algorithm provided in Eq. (16), where the coefficient updates can be accomplished by solving a linear system while the partition can be updated using a stochastic gradient method.
@@ IV Numerical Studies
This work utilized two numerical studies to assess the proposed NF representation and proposed ProxNF method for spatiotemporal image reconstruction. These studies were inspired by dynamic contrast-enhanced photoacoustic computed tomography (DCE PACT) imaging of tumor perfusion in small animal models [52]. PACT is a hybrid biomedical imaging modality that combines benefits from both optical and ultrasound imaging[53]. Specifically, PACT combines the benefits of high soft tissue contrast from optical imaging and high-resolution detection principles from ultrasound imaging to create high-resolution maps of optical properties. PACT image formation utilizes a short laser pulse in the near infrared range to illuminate a biological tissue. The biological tissue then absorbs the optical energy and generates heat and a local increase in pressure, which is described as the photoacoustic effect. This pressure distribution then propagates as an acoustic wave where it can be measured outside the tissue using ultrasound detectors. Under the assumption of a non-lossy homogeneous acoustic medium, that is, a medium with homogeneous speed of sound and density and no acoustic attenuation, the relationship between the induced pressure distribution and the data can be modeled using the circular Radon transform in two spatial dimensions and the spherical Radon transform in three spatial dimensions [54, 55]. The numerical studies presented in this work, focus on estimation of this spatially and temporally varying induced pressure distribution from sparse-view measurements collected from a circular aperture in two spatial dimensions and a cylindrical aperture in three spatial dimensions.
@@ IV-A Construction of the dynamic object
The dynamic objects used in these studies were spatiotemporal maps of the spatially and temporally varying induced pressure distribution within a numerical tumor perfusion model in two and three spatial dimensions. The anatomical structures of these objects were constructed utilizing an abdominal section extracted from an anatomically realistic whole-body murine numerical phantom (MOBY) [56, 57]. MOBY comprises detailed three-dimensional (3D) anatomical structures featuring tens of organs and models physiologically realistic cardiac and respiratory motions and is illustrated in Figure 1. A spherical lesion of diameter 7 mm was inserted in the abdominal region at a depth of 5 mm to mimic a subcutaneous tumor flank. To simulate DCE imaging of a tumor perfusion model, the concentration of an optical contrast agent in the lesion and blood-filled organs was computed using a two compartment Toft’s model [58, 59, 60]. In this model, the contrast agent was introduced after 90 seconds and injected over a 30 second window following a slow bolus pattern. Optical properties (optical absorption coefficient and scattering coefficient) corresponding to an 800 nm laser pulse wavelength were then assigned to each tissue type based on physiologically plausible chromophore concentrations (oxygenated hemoglobin, deoxygenated hemoglobin, water, fat, contrast agent)[61]. Contrast agent optical properties were chosen to mimic those of indocyanine green (ICG) [62], in which case the absorption spectrum has a peak near the chosen wavelength.
The diffusion approximation[63] was used to simulate photon transport at each imaging frame. The incident fluence was assumed constant on the surface of the object. The fluence distribution within the object was obtained by solving the partial differential equation stemming from the diffusion approximation using the open-source finite element library FEniCS [64], thus yielding a volumetric map of dynamic induced pressure distribution at each imaging frame. For the experiments in two spatial dimensions, a slice was extracted from the 3D induced pressure phantom.
Figure 1: Numerical dynamic mouse abdominal phantom extract from mouse whole body phantom (MOBY) and shown at the beginning and end of inspiration during a 5 second breathing period. The phantom is comprised of detailed three-dimensional (3D) anatomical structures featuring several organs and models physiologically realistic cardiac and respiratory motions. A spherical lesion of diameter 7 mm was inserted in the abdominal region at a depth of 5 mm to mimic a subcutaneous tumor flank and displayed in dark red.
@@ IV-B1 Dynamic imaging system for two spatial dimensions
Figure 2 illustrates the virtual imaging system used in the numerical study with two spatial dimensions. The measurement geometry was a circle of radius $R$ (shown in green), along which two groups of 5 idealized point-like ultrasound transducers are distributed. The sensors in each group were spaced $1^{\circ}$ apart on the transducer arc. Each group of sensors was distributed antipodal from one another. The object was placed at the center of the imaging system and fully contained within a square region of size $L$ (field of view). At each imaging frame, data were collected simultaneously by each sensor corresponding to the integral along $I$ concentric arcs. Sensors were rotated by an angle $\Delta\Theta$ after each imaging frame. The dynamic changes of the object were assumed to be slow enough such that the object was not expected to vary significantly during the acquisition of each imaging frame (quasi-static approximation). This assumption is often valid for several dynamic PACT applications since the duration of the light excitation pulse at each imaging frame is in the range of a few nanoseconds while the dynamic changes in the object have a time scale in the range of milliseconds to seconds.
Under these assumptions, the imaging operator in two spatial dimensions can then be described by the circular Radon transform (CRT). The discrete-to-discrete counterpart of the CRT imaging operator was implemented using the AirToolsII MatLab toolbox [65] and stored as a compressed sparse column matrix[66]. The imaging systems parameters are displayed in Table I.
Figure 2: Illustration of the virtual system for dynamic imaging using a CRT model. At each imaging frame, data are collected by idealized point-like transducers (shown in blue). Transducers rotate an angle of $\Delta\Theta$ between frames.
@@ IV-B2 Dynamic imaging system for three spatial dimensions
The object was placed at the center of the imaging system and fully contained within a rectangular prism region of side length $L$ in the $x$ and $y$ direction and height $H$. At each imaging frame, data were collected at each sensor ($I$ measurements per sensor). Sensors rotated an angle $\Delta\Theta$ along the z-axis of the cylindrical geometry at each frame. Again, the dynamic changes in the object were assumed to be slow enough such that the object is approximately constant at each imaging frame.
@@ IV-C Study Design
The first numerical study solved the embedding problem and assessed the proposed NF architecture for representing the dynamic object described in Section IV-A. For the case involving two spatial dimensions, the proposed NF architecture was compared to the original NF architecture with polynomial basis functions [13] and a low rank (semiseparable) decomposition of the dynamic object. The proposed POUnet architecture used a network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ with 4 hidden layers of width 140 nodes and one output layer of width $P=10$ partitions. Sinusoidal and softmax activation functions were employed in the hidden and output layers, respectively. The set of basis function $\mathbf{B}$ consisted of $L=186^{2}=34596$ pixel indicator functions. This resulted in a representation with $427,020$ parameters. The original POUnet architecture used a network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ with 5 hidden layers of width 280 nodes and an output layer of width $P=100$ partitions. Third order space-time separable polynomials were employed as basis functions ($L=40$). This resulted in a representation with $425,800$ parameters. These two NF representations were compared to a semiseparable approximation of rank $r=12$, SS(12), which has $430,704$ parameters. These representations allow for a 40-fold reduction in the number of parameters compared to representing the object as a collection of spatiotemporal voxel values ($\sim 20$ million parameters).
For the case involving three spatial dimensions, the proposed NF architecture was compared to a low rank, semiseparable, decomposition of the dynamic image. The POUnet architecture again utilized the same network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ as the 2D case and a set of basis function $\mathbf{B}$ that contained $L=186^{2}\cdot 100=3,459,600$ voxel indicator functions. This resulted in a representation with $34,988,424$ parameters. The semiseparable approximation used a rank $r=10$, SS(10), which has $34,608,960$ parameters. These representations yield over two orders of magnitude reductions in the number of parameters compared to representing the object as a collection of spatiotemporal voxel values ($\sim 5$ billion parameters).
@@
The second numerical study explored dynamic image reconstruction in two (CRT imaging operator) and three (SRT imaging operator) spatial dimensions. In both cases, two spatiotemporal image reconstruction approaches were utilized. The first reconstruction technique utilized the proposed NF architecture and the proposed ProxNF method in Algorithm 1. The second reconstruction technique, which serves as a reference, was spatiotemporal image reconstruction with nuclear norm regularization (STIR-NN) [69]. In particular, for the case involving three space dimensions, the memory-efficient maximal-rank-constrained method in [11] was implemented. For the two spatial dimensions case, the NF approach in [13], which directly minimizes the objective in Eq. (14) using gradient descent, was also considered.
Measurement data, virtually acquired as described in Sect. IV-B, were corrupted with i.i.d. additive Gaussian noise with zero mean and standard deviation $\sigma$ controlled by a specified relative noise level (RNL), defined as $\operatorname{RNL}=\sigma/\|\boldsymbol{d}\|_{\infty}$. In the two spatial dimensions case, the RNL was set to 0.04 and, in the three spatial dimensions case, the RNL was set to 0.01. Regularization weights for all reconstructions were chosen using Morozov’s discrepancy principle [70], which selects the regularization weight such that the optimal data-fidelity value is approximately equal to the noise standard deviation.
@@
In these studies, image quality was quantified using relative root mean square error (RRMSE) and structural similarity index measure (SSIM) [71], evaluated both on the entire image and on a region of interest (ROI) containing the lesion. Image quality was also assessed by analyzing the estimated lesion activity curve (LAC), which is the time-varying spatially-averaged signal over a volume of interest (the lesion).
@@ V Results
The proposed method, ProxNF, was implemented using PyTorch, an open-source machine learning framework [72] and executed on GPUs. The POUnet parameters $\boldsymbol{\eta}$ and $\boldsymbol{C}$ were initialized by performing an embedding of the time-averaged object for the representation studies and by solving a static image reconstruction problem on time-averaged measurements for the reconstruction studies.
Following initialization, the NFs in the representation study were trained using the alternating minimization approach in Eq. (16). Similarly, after initialization, the NFs in the reconstruction study were trained with the proposed proximal splitting method in Algorithm 1. A reduction of two orders of magnitude in the norm of the proximal gradient was used as a stopping criterion. The batch size was $J=324$ and time frames were stochastically selected at each iteration. Tikhonov regularization was used to penalize the $L^{2}$ norm of the gradient of the spatiotemporal image. The regularized embedding problem at each proximal update was solved as follows. The linear systems corresponding to the updates of $\boldsymbol{C}$ were solved using the conjugate gradient method. The internal updates for $\boldsymbol{\eta}$ were accomplished using an Adam optimizer [24]. Specifically, 10,000 iterations with a fixed learning rate of $10^{-5}$ and batch size of 100,000 randomly sampled spatiotemporal points were used.
In the two-dimensional case, an additional NF was trained utilizing the method in [13], referred in what follows as GradNF. GradNF directly minimizes the objective in (14) using gradient descent and the alternating minimization algorithm in Eq. (16). For ease of comparison, GradNF utilized the same network architecture and randomization approach as the ProxNF method. The internal updates of $\boldsymbol{C}$ and $\boldsymbol{\eta}$ were similarly accomplished using an Adam optimizer [24] with a fixed learning rate of $10^{-5}$ and batch sizes of $1.1.\times 10^{7}$ corresponding to spatiotemporal voxels for each imaging frame. A reduction of two orders of magnitude in the norm of the gradient was used as a stopping criterion.
The spatiotemporal reconstructions in two spatial dimensions using the reference method (STIR-NN) were accomplished utilizing the Fast Iterative Shrinkage Algorithm (FISTA) [73] as implemented in UNLocBoX, a MATLAB toolbox for proximal-splitting methods [74]. The low rank semiseparable approximation of the image with two spatial dimensions was accomplished using the SVD command in MATLAB and soft-thresholding. The spatiotemporal reconstructions in three spatial dimensions using the reference method (STIR-NN) were accomplished using the memory-efficient randomized algorithm in [11] with a maximum rank constraint of 32. The randomized singular value decomposition [75] was employed to compute the low rank decomposition and enforce the maximum rank constraint.
@@ V-A Study 1: Dynamic embedding problem
Figure 3: Study 1a: Dynamic embedding in two spatial dimensions. Lesion activity curve (LAC) of the object (blue), SS(12) approximation (green), original POUnet architecture (purple), and proposed POUnet architecture (red). The proposed POUnet architecture better captures the dynamic changes of the object compared to the original POUnet architecture in [13].
Figure 4: Study 1a: Dynamic embedding in two spatial dimensions. Object (top), Rank 12 semiseparable approximation SS($12$) (second row), NF using original POUnet architecture (third row), and NF using proposed POUnet architecture are shown at selected frames throughout the time period. The dynamic range of the grayscale image is $[10^{-}4,0.06]$. A logarithmic scale was used to increase visualization contrast at depth. An animation of each representation constructed in this study is available in the supplemental multimedia materials (Video 1).
Figure 5: Study 1b: Dynamic embedding in three spatial dimensions. Lesion activity curve (LAC) of the object (blue), the semi-separable rank 10 approximation (green), and the NF representation (red). The semiseparable and NF representations lead to almost identical, smooth LACs that underestimates activity at later times.
Figure 6: Study 1b: Dynamic embedding in three spatial dimensions. Object (top), Rank 10 semiseparable approximation SS($10$) (second row), NF representation (bottom row) with proposed POUnet architecture are shown at selected frames throughout the time period. ParaView [76] was used for volume rendering. Dynamic range of the rendering is [0, 0.06] using a linear scale. Both methods provide high fidelity representations but slightly underestimate the tumor activity in later time periods. An animation of each representation constructed in this study is available in the supplemental multimedia materials (Video 2).
TABLE II: Study 1: Dynamic embedding. RRMSE, SSIM, ROI RRMSE, ROI SSIM, and LAC RRMSE achieved by the NF representations with original and proposed POUnet architectures and the rank $r$ semiseparable approximations SS($r$) in two (2D) and three (3D) spatial dimensions.
@@
The accuracy metrics—RRMSE and SSIM over the entire image, RRMSE and SSIM over the region of interest containing the inserted lesion, and the lesion activity curve (LAC) RRMSE—achieved by each representation are reported in Table II. The LACs associated with these representations and with the object are shown in Fig. 3. Selected frames of these representations and the object are shown in Fig. 4. The proposed NF architecture showed superior approximation properties (as quantified by the above accuracy metrics) compared to the original POUnet architecture in [13], and comparable accuracy to the semiseparable approximation with a similar number of parameters.
@@
The accuracy metrics for the proposed POUnet representation and the semiseparable approximation of rank 10 (SS(10)) are reported in Table II. The LACs associated with these representations and with the object are shown in Fig. 5. Selected frames of these representations and the object are shown in Fig. 6. The proposed NF architecture achieved comparable accuracy metrics to those achieved by the semiseparable approximation with a similar number of parameters.